# How Conflict Aversion Can Enable Authoritarianism: An Evolutionary Dynamics Approach

### v1

@@ Abstract
We use evolutionary game theory to examine how conflict-averse centrism can unintentionally facilitate authoritarian success in polarized political conflicts. Many such conflicts are asymmetric: authoritarian actors can employ norm-breaking or coercive tactics, while democratic resistance faces stronger constraints on what counts as normatively acceptable behavior. Yet formal models typically treat opposing sides symmetrically and rarely examine conflict-averse behavior. Drawing on empirical research on protest backlash, civility norms, and authoritarian resilience, we model these dynamics as a three-strategy evolutionary game in which resistance, authoritarianism, and conflict-averse centrism interact under replicator dynamics. This framework yields two distinct outcomes—cyclic resurgence of authoritarian strength through a heteroclinic cycle and a stable centrist–authoritarian coalition that excludes resistance—depending on how actors respond to confrontation. The analysis shows how payoff differences can reorganize long-run dynamics in asymmetric conflicts. Our contribution is to demonstrate how an established dynamical framework, combined with empirically grounded behavioral assumptions, clarifies the strategic conditions under which conflict aversion can diminish the effectiveness of democratic resistance.
@@ 1 Introduction
Although such actors often remain muted or invisible in media discourse (Bail2021), they constitute a large share of the mass public (Mutz2018) and shape the strategic landscape in which opposing forces interact. Despite this empirical importance, conflict aversion rarely appears as a formalized behavioral strategy in theoretical models. Most models of polarization treat individuals as points in a continuous metric space and define polarization as a distributional property of those points (FriedkinJohnsen1999; Deffuant2000; HegselmannKrause2002; FlacheMacy2011; TurnerSmaldino2018; Duggins2017; FlacheEtAl2017). These frameworks model contending poles symmetrically: their influence equations do not distinguish differences in normative content or strategic objectives. Related evolutionary models that include centrist or moderate strategies likewise interpret centrism as an ideological midpoint rather than a preference for minimizing confrontation (e.g., ArceSandler2003; ShortEtAl2017).
These patterns suggest that conflict aversion functions as a distinct strategic type with its own payoff structure, motivating a game-theoretic approach. We therefore develop a minimal evolutionary-game model with three behavioral strategies: authoritarianism or fascism, resistance to authoritarian encroachment, and conflict-averse centrism that imposes reputational or social penalties on visible confrontation. We encode strategy payoffs in a three-by-three game and study their population dynamics using replicator dynamics (TaylorJonker1978; HofbauerSigmund1998).
We then analyze two empirically motivated strategic regimes: a cyclic resurgence regime in which resistance locally outperforms authoritarianism but conflict-averse centrism generates a heteroclinic cycle, and a coalition regime in which mutually reinforcing centrists and authoritarians stabilize against resistance. Across these regimes, empirically grounded assumptions about resistance, conflict aversion, and authoritarian norm-breaking overturn the expectation that local dominance of resistance over authoritarianism guarantees long-run success.
@@ 2 Background
These frameworks share key structural features: opinions are represented as points in a metric space, and polarization is a distributional property of those points. Interacting poles are modeled symmetrically, and the equations governing influence do not distinguish normative content or strategic objectives. Even when heterogeneity in responsiveness is included, the models do not treat behavior shaped primarily by conflict aversion as a distinct type. Thus, they illuminate how clustering emerges from local influence rules but offer limited insight into how conflict-averse behavior might alter asymmetric strategic contests. Evolutionary models with multiple ideological or attitudinal positions likewise treat “moderate” types as ideologically intermediate rather than as actors motivated by minimizing confrontation (e.g., ArceSandler2003; ShortEtAl2017).
@@ 3 Modeling
We consider a population composed of three behavioral strategies: resistance ($R$), authoritarian fascism ($F$), and conflict-averse centrism ($C$). Let $x(t)=(x_{R},x_{C},x_{F})$ denote the population frequencies of these strategies at time $t$, constrained to the simplex
@@ 4 Results
where $D>0$ is the normalizing constant that arises when expressing the equilibrium conditions in coordinates. This expression follows from standard linearization formulas for interior equilibria in three-strategy replicator systems (HofbauerSigmund1998). When
@@
then we say the heteroclinic cycle is $F$-dominated: along the cycle, trajectories spend an asymptotically larger fraction of each successive excursion in neighborhoods of $e_{F}$ than near the other vertices, so sufficiently long empirical time averages can be made arbitrarily close to the fascist vertex. This outcome is generated entirely by the payoff sign structure: although resistance defeats fascism in isolation, the presence of conflict-averse centrists produces a global cycle in which fascism dominates long stretches of the trajectory.
@@
Here $b_{R}$ measures resistance’s advantage against fascists; $a_{R}$ captures the penalty centrists impose on resistance; and $c_{C}$ and $c_{F}$ represent the mutual benefits within the CCFF coalition. Notably, increasing $c_{C}$ tightens the anti-invasion condition, making invasion harder, whereas increasing $c_{F}$ relaxes it, making invasion easier.
@@
Under the same condition, any interior rest point of the replicator dynamics would require all strategies to earn equal payoff, which implies the necessary condition
@@ 5 Discussion and Conclusions
The first mechanism shows that even when resistance is intrinsically stronger than authoritarianism in direct confrontation, the presence of conflict-averse actors can generate a heteroclinic cycle. In parameter regions where the eigenvalue ratios favor fascism, this cycle produces a regime in which fascism repeatedly resurges and can dominate long stretches of the trajectory. In those same regions—where trajectories spend disproportionately long near the fascist vertex—this structure offers a formal account of how civility norms, media framing of protest as disruptive, and public discomfort with confrontation can repeatedly regenerate authoritarian strength. The resulting cycling implies that democratic movements may experience recurring episodes in which authoritarian actors come close to dominating the population, even though resistance would eliminate authoritarianism were centrists absent. Conflict aversion thus does more than slow democratic victory: it reorganizes the global dynamics in ways that prevent stable democratic dominance.
The second mechanism reveals a distinct pathway through which conflict aversion can advantage authoritarian projects. When centrists impose penalties on resistance that are sufficiently strong—relative to resistance’s advantage against fascism and to the coalition parameters $c_{C}$ and $c_{F}$ that govern mutual benefits within the centrist–fascist coalition, as captured by the invasion condition $b_{R}<c_{C}(1+a_{R}/c_{F})$—the two may form a dynamically stable coalition along the CCFF edge that resistance cannot invade. Notably, increasing $c_{C}$ tightens this anti-invasion condition by raising the right-hand side, whereas increasing $c_{F}$ relaxes it by appearing in the denominator. Although resistance still defeats fascism in isolation, the alliance between conflict-averse actors and authoritarian actors can lock the system into a boundary equilibrium with a persistent and potentially high fascist share. The coalition’s ability to exclude resistance depends jointly on the cost $a_{R}$ that centrists impose on resistance and the mutual reinforcement parameters $c_{C}$ and $c_{F}$; these same parameters determine how close the equilibrium lies to the fascist vertex. This mechanism formalizes how appeals to civility, moderation, or both-sides norms can unintentionally sustain authoritarian projects by generating a stable alignment that structurally disadvantages democratic resistance.

### v2

@@ Abstract
We use evolutionary game theory to examine how conflict-averse centrism can facilitate authoritarian success in polarized political conflicts. Such conflicts are often asymmetric: authoritarian actors can employ norm-breaking or coercive tactics, while democratic resistance faces stronger normative constraints on acceptable behavior. Yet formal models typically treat sides symmetrically and rarely examine conflict-averse behavior. Drawing on empirical research on protest backlash, civility norms, and authoritarian resilience, we model these dynamics as a three-strategy evolutionary game. This framework yields two outcomes—cyclic authoritarian resurgence through a heteroclinic cycle and a stable centrist–authoritarian coalition excluding resistance—depending on confrontation responses. We demonstrate how an established dynamical framework with empirically grounded behavioral assumptions clarifies conditions under which conflict aversion can diminish the effectiveness of democratic resistance.
@@ 1 Introduction
Although such actors often remain muted or invisible in media discourse (Bail2021), they constitute a large share of the mass public (Mutz2018) and shape the strategic landscape in which opposing forces interact. Despite this empirical importance, conflict aversion rarely appears as a formalized behavioral strategy in theoretical models. Most models of polarization treat individuals as points in a continuous metric space and define polarization as a distributional property of those points (FriedkinJohnsen1999; Deffuant2000; HegselmannKrause2002; FlacheMacy2011; TurnerSmaldino2018; Duggins2017; FlacheEtAl2017). These frameworks model contending poles symmetrically with respect to normative content and strategic objectives: their influence equations do not distinguish differences in normative content or strategic aims. Related evolutionary models that include centrist or moderate strategies likewise interpret centrism as an ideological midpoint rather than a preference for minimizing confrontation (e.g., ArceSandler2003; ShortEtAl2017).
These patterns suggest that conflict aversion functions as a distinct strategic type with its own payoff structure, motivating a game-theoretic approach that treats centrists’ responses to resistance and authoritarianism asymmetrically. We therefore develop a minimal evolutionary-game model with three behavioral strategies: authoritarianism or fascism, resistance to authoritarian encroachment, and conflict-averse centrism that imposes reputational or social penalties on visible confrontation. We encode strategy payoffs in a three-by-three game and study their population dynamics using replicator dynamics (TaylorJonker1978; HofbauerSigmund1998).
We then analyze two empirically motivated strategic regimes: a cyclic resurgence regime in which resistance defeats authoritarianism in isolation but conflict-averse centrism generates a heteroclinic cycle, and a coalition regime in which mutually reinforcing centrists and authoritarians stabilize against resistance. Across these regimes, empirically grounded assumptions about resistance, conflict aversion, and authoritarian norm-breaking overturn the expectation that pairwise dominance of resistance over authoritarianism guarantees long-run success.
@@ 2 Background
These frameworks share key structural features: opinions are represented as points in a metric space, and polarization is a distributional property of those points. Interacting poles are modeled symmetrically in the sense that the equations governing influence do not distinguish normative content or strategic objectives. Even when heterogeneity in responsiveness is included, the models do not treat behavior shaped primarily by conflict aversion as a distinct type. Thus, they illuminate how clustering emerges from local influence rules but offer limited insight into how conflict-averse behavior might alter asymmetric strategic contests. Evolutionary models with multiple ideological or attitudinal positions likewise treat “moderate” types as ideologically intermediate rather than as actors motivated by minimizing confrontation (e.g., ArceSandler2003; ShortEtAl2017).
@@ 3 Modeling
We consider a population composed of three behavioral strategies: resistance ($R$), authoritarian fascism ($F$, “fascism” for short), and conflict-averse centrism ($C$). Let $x(t)=(x_{R},x_{C},x_{F})$ denote the population frequencies of these strategies at time $t$, constrained to the simplex
@@
For the RRFF interaction, resistance gains a payoff advantage over fascism ($A_{RF}=b_{R}>0$) while fascism incurs a penalty ($A_{FR}=-a_{F}<0$), consistent with $R\succ F$. For the CCFF interaction, the two regimes we study assume either that fascism exploits centrists ($A_{CF}=-a_{C}<0$ and $A_{FC}=b_{F}>0$, yielding $F\succ C$) or that $C$ and $F$ benefit simultaneously ($A_{CF}=c_{C}>0$ and $A_{FC}=c_{F}>0$, yielding $C\leftrightarrow F$).
@@ 4 Results
where $D>0$ is a positive polynomial in the payoff parameters that arises when solving (19) for $x^{\ast}$ and substituting into the Jacobian; its explicit form is not required here because only its sign matters, and positivity follows from the strict cyclic-dominance structure of $A^{(1)}$. This expression follows from standard linearization formulas for interior equilibria in three-strategy replicator systems (HofbauerSigmund1998). When
@@
then we say the heteroclinic cycle is $F$-dominated: along the cycle, trajectories spend an asymptotically larger fraction of each successive excursion in neighborhoods of $e_{F}$ than near the other vertices, so sufficiently long empirical time averages can be made arbitrarily close to the fascist vertex, even though the trajectory itself does not converge to that vertex. This outcome is generated entirely by the payoff sign structure: although resistance defeats fascism in isolation, the presence of conflict-averse centrists produces a global cycle in which fascism dominates long stretches of the trajectory.
@@
Here $b_{R}$ measures resistance’s advantage against fascists; $a_{R}$ captures the penalty centrists impose on resistance; and $c_{C}$ and $c_{F}$ represent the mutual benefits within the CCFF coalition. Notably, increasing $c_{C}$ enlarges the parameter region in which resistance cannot invade, making invasion harder, whereas increasing $c_{F}$ shrinks that region, making invasion easier.
@@
Any interior rest point of the replicator dynamics would require all strategies to earn equal payoff; in particular, setting $\pi_{R}=\pi_{C}$ yields the necessary condition
@@ 5 Discussion and Conclusions
The first mechanism shows that even when resistance is intrinsically stronger than authoritarianism in direct confrontation, the presence of conflict-averse actors can generate a heteroclinic cycle. In parameter regions where the eigenvalue ratios favor fascism, this cycle produces a regime in which fascism repeatedly resurges and can dominate long stretches of the trajectory. In those same regions—where trajectories spend disproportionately long near the fascist vertex—this structure offers a formal account of how civility norms, media framing of protest as disruptive, and public discomfort with confrontation can repeatedly regenerate authoritarian strength. The resulting cycling implies that democratic movements may experience recurring episodes in which authoritarian actors come close to dominating the population, even though resistance would eliminate authoritarianism were centrists absent. Conflict aversion thus does more than impede resistance: it reorganizes the global dynamics in ways that prevent stable democratic dominance.
The second mechanism reveals a distinct pathway through which conflict aversion can advantage authoritarian projects. When centrists impose penalties on resistance that are sufficiently strong—relative to resistance’s advantage against fascism and to the coalition parameters $c_{C}$ and $c_{F}$ that govern mutual benefits within the centrist–fascist coalition, as captured by the invasion condition $b_{R}<c_{C}(1+a_{R}/c_{F})$—the two may form a dynamically stable coalition along the CCFF edge that resistance cannot invade. Notably, increasing $c_{C}$ raises the right-hand side of this inequality, expanding the parameter region in which invasion fails, whereas increasing $c_{F}$ lowers it, shrinking that region. Although resistance still defeats fascism in isolation, the alliance between conflict-averse actors and authoritarian actors can lock the system into a boundary equilibrium with a persistent and potentially high fascist share. The coalition’s ability to exclude resistance depends jointly on the cost $a_{R}$ that centrists impose on resistance, resistance’s advantage $b_{R}$ against fascists, and the mutual reinforcement parameters $c_{C}$ and $c_{F}$, while the equilibrium’s location along the CCFF edge—and thus how close it lies to the fascist vertex—is determined solely by $c_{C}$ and $c_{F}$. This mechanism formalizes how appeals to civility, moderation, or both-sides norms can unintentionally sustain authoritarian projects by generating a stable alignment that structurally disadvantages democratic resistance.
@@
Acknowledgments
The author used Chat-GPT 5.1 to refine and tighten text and to proofread this manuscript.